OpenStudy (anonymous):

the length of a hypotenuse if one right angle is 7 and the other is 9 how do you find the length

OpenStudy (anonymous):

Pythagorean theorem. a^2+b^2=c^2. let a=7, b=9, solve for c. 7^2+9^2=c^2 49+81=c^2 c^2=130 c=sqrt(130)

OpenStudy (anonymous):

\[h^2 = a^2 + b^2\] so \[h^2 = 7^2 +9^2\] \[h^2 = 49 + 81\] \[h^2 = 130\] \[h = \sqrt{130}\]

OpenStudy (anonymous):

thanks for helping out!